OpenStudy (anonymous):

The sum of two numbers is 110. Their difference is 28. What is the smallest of the two numbers?

OpenStudy (anonymous):

A. 56 B. 41 C. 69 D. 55

OpenStudy (anonymous):

x + y = 110 x - y = 28. what do you think? assume y<x

OpenStudy (anonymous):

btw. x = 69,y=41. but don't look at the answer. find out why :) first.!

OpenStudy (anonymous):

Well, it says which is the smallest of the two numbers, so I'd have to go with 41 because that is smaller.

OpenStudy (anonymous):

Is tha

OpenStudy (anonymous):

Is that right or am I wrong about that one?

OpenStudy (anonymous):

41+69=110 looks good to me

OpenStudy (mathmate):

Here's the trick: bigger number = (sum+difference)/2 smaller number = (sum-difference)/2

OpenStudy (anonymous):

I am so confused on which one it is. :/

OpenStudy (anonymous):

So it must be 41 because that is the smaller number.

OpenStudy (mathmate):

If you want the smaller number, Calculate (110-28)/2

OpenStudy (anonymous):

I calculated that and got 96.

OpenStudy (anonymous):

@alain relax you have it!

OpenStudy (anonymous):

you have the answer. use it :P

OpenStudy (anonymous):

Why does everyone call me Alain? Lol.. my name is Alaina. xD Oh well, it can be short for it. :P And thanks everyone! :)

OpenStudy (anonymous):

there only is one solution.

OpenStudy (anonymous):

MMMKKKKK!!!!

OpenStudy (anonymous):

So 41 right?

OpenStudy (anonymous):

oh and btw. (110-28) /2 = 41 :P

OpenStudy (anonymous):

check your subtraction ^^

OpenStudy (anonymous):

RONNCC has a good answer, basically it is best to use substitution in order to solve. He said: "x + y = 110 x - y = 28. what do you think? assume y<x" and you may assume without loss of generality that y<x, since x-y=28 > 0 meaning x>y. y=x-28 from the second equation, so just substitute this into the first to give: x+x-28 = 110 ---> 2x=138 --> x=69. To find y simply use your first equation: 69+y=110 ---> y=41.

OpenStudy (anonymous):

substitution is nice for harder equations, but generally elimination is a good method when the coefficient of x is 1 and they are in a simple form like this.